while hovering near the top of a waterfall in national park at feet, a helicopter pilot accidentally drops his sunglasses. The height h(t) of sunglasses after t seconds is given by polynomial function h(t)=-16t^(2)+1936. when will the sunglasses hit the ground?
Answer by Cromlix(4381)

h(t)=-16t^(2)+1936.
When glasses hit ground h(t) = 0
0= -16t^(2)+1936.
- 16t^2 = -1936
16t^2 = 1936
t^2 = 1936/16
t^2 = 121
t = 11
His glasses will take 11 seconds to hit the ground.
Hope this helps.
:-)